OpenStudy (anonymous):

The cost of manufacturing x cases of fillfluff is C dollars (in thousands), where C(x) = 4 x + 4 sqrt(x) + 5 Monthly production at t weeks is expected to be x = 6400 + 90 t Write an equation to express rate of change of cost with respect to time: (dC)/(dt) =

OpenStudy (anonymous):

\[\frac{dC}{dt}=\frac{dC}{dx}\frac{dx}{dt}\]

OpenStudy (anonymous):

\[\frac{dx}{dt}=90\] \[\frac{dC}{dx}=4+\frac{2}{\sqrt{x}}\]

OpenStudy (anonymous):

thats the marginal cost

OpenStudy (anonymous):

how can you find the rate of change?